OpenStudy (anonymous):

HELP PLEASE. a youth baseball league recently completed registration. the league was able to form 15 team of 11 players each how many players registered?

OpenStudy (rosedewittbukater):

Isn't it just 15x11? 165 I think if it is.

OpenStudy (mathstudent55):

1 team has 11 = 1 x 11 = 11 players 2 teams have 11 + 11 = 2 x 11 = 22 players 3 teams have 11 + 11 + 11 = 3 x 11= 33 players 4 teams have 11 + 11 + 11 + 11 = 4 x 11 = 44 players ... 15 teams have ? x 11 = ?? players Do you see the pattern there? This is a multiplication problem. What numbers go in ? and ??

OpenStudy (anonymous):

ya i see and thanks i get it

OpenStudy (anonymous):

great method of telling mathstudent55

OpenStudy (anonymous):

what do you mean

OpenStudy (anonymous):

I was praising his method of explanation :)

OpenStudy (mathstudent55):

Thanks shannycampos & AtiFS